The diameters of ball bearings are distributed normally. The mean diameter is 123123 millimeters and the standard deviation is 22 millimeters. Find the probability that the diameter of a selected bearing is greater than 125125 millimeters. Round your answer to four decimal places.
Answer #1

Solution :

Given that ,

mean = = 123

standard deviation = = 2

P(x > 125) = 1 - P(x < 125)

= 1 - P[(x - ) / < (125 - 123) / 2)

= 1 - P(z < 1)

= 1 - 0.8413

= 0.1587

Probability = 0.1587